# ConsentDiff at Scale: Longitudinal Audits of Web Privacy Policy Changes and UI Frictions

### v2

@@ Abstract.
Web privacy is experienced via two public artifacts: site utterances in policy texts, and the actions users are required to take during consent interfaces. In the extensive cross-section audits we’ve studied, there is a lack of longitudinal data detailing how these artifacts are changing together, and if interfaces are actually doing what they promise in policy. ConsentDiff provides that longitudinal view. We build a reproducible pipeline that snapshots sites every month, semantically aligns policy clauses to track clause-level churn, and classifies consent-UI patterns by pulling together DOM signals with cues provided by screenshots. We introduce a novel weighted claim–UI alignment score, connecting common policy claims to observable predicates, and enabling comparisons over time, regions, and verticals. Our measurements suggest continued policy churn, systematic changes to eliminate a higher-friction banner design, and significantly higher alignment where rejecting is visible and lower friction.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Human and societal aspects of security and privacyccs: Security and privacy Usability in security and privacyccs: Information systems Web miningccs: Information systems World Wide Web
@@ 1. Introduction
Online privacy is enacted through two public-facing artefacts: (i) privacy policies, which disclose purposes, retention, and sharing, and (ii) consent interfaces (CMP banners/modals), which implement user choice. Prior research observe dishonest trends and pervasive violations in cookie banners and consent flows (nouwens2020cookiebanners; matte2020cookierespect; mathur2019darkpatterns), but a longitudinal,web-scale view that jointly tracks what sites say (policy text) and what users must do (consent UI patterns) and a concrete metric for whether these align.
This is a problem for both end users and regulators. Policies are usually modified in the absence of any notice—be it through a template refresh, a vendor switch, or due to edits from a given jurisdiction—while banners change through A/B testing and the default settings of CMPs. Audits at a single point in time will pick up on the most blatant violations, but they miss how deployments change over time (e.g., several months), which patterns persist, and whether textual claims are keeping up with changes in the UI. On the other hand, human consumable guidance and vendor frameworks (e.g., how a compliant consent process would look (edpb2020consent; iab2020tcf)) lay out what compliant consent should look like but do not offer a reproducible way of measuring the degree to which claims (e.g., “opt-in,” “easy reject”) map onto action steps users take.
We present ConsentDiff, a measurement framework that jointly tracks policy text and consent interfaces over time. Each month, we snapshot a stratified set of sites, semantically align policy clauses across versions, classify consent-UI patterns from DOM and screenshots, and compute a claim–UI alignment score that pairs policy statements with observable defaults and actions. Our work complements research on temporal drift in privacy recall which models how people misremember original sharing audience (guo2025temporaldriftprivacyrecall); our focus is on the site-side artefacts, as well as noting their evolution.
This paper contributes:
  1. (1)
    A comprehensive, replicable auditing framework that connects changes to policies at the clause-level to a large-scale, monthly consent UI classified by policies, via semantic and edit alignment of policy clauses, and a weakly supervised model of DOM+vision for interface patterns. We describe performance and failure modes.
  2. (2)
    An interpretable metric for claim-UI alignment that formalizes and evaluates mappings from policy claim (e.g., opt-in, minimization) to observable UI predicates using weights, sensitivity analysis, and ablations.
  3. (3)
    A longitudinal measurement study that captures and quantifies temporal trends, shifts in patterns, cross-region/vertical gaps, and an event response, complete with effect sizes, confidence intervals, fixed-effects estimates, and robustness checks.
@@ 2.1. Consent Interfaces and Dark Patterns
Large-scale audits have documented pervasive lack of compliance with and the use of dark-pattern designs in consent banners and CMP deployments, which illustrate how choices in interface can nudge users towards acceptance and obscure paths to rejection (nouwens2020cookiebanners; matte2020cookierespect; mathur2019darkpatterns; degeling2019wevalue). Practice and policy guidance add further specificity to valid consent elements (e.g., unambiguous opt-in, symmetry of choices, easy withdrawal) and have codified vendor-facing frameworks that many sites adopt (edpb2020consent; iab2020tcf). Outside of consent dialogs, dark-pattern taxonomies and crawls have captured patterns of manipulative designs across e-commerce and platforms, including intrusion, default bias, and interference with interfaces (mathur2019darkpatterns; gray2021). Our work connects these dimensions by tracking both the text a site publishes and the UI it offers over time, then quantifying their consistency with a claim–UI alignment score.
@@ 2.2. Automated Policy Analysis and Web-Scale Measurement
Automated tools segment and label privacy policies to enable designed query and compliance studies on a massive scale (harkous2018polisis; lippi2019claudette). Extending this area of work, we align policy clauses over time to measure clause-level churn, and to relate this churn to consent UI changes. There are research-oriented rankings that utilize methodology to reduce ”bias” in the top lists for our sampling and instrumentation (pochat2019tranco), which is supported with classic web-tracking studies on the importance of maintaining long term request/cookie logs and distinguishing between first/third parties, capturing screenshots indicating the UI states during data collection methods (englehardt2016online; acar2014webneverforgets; lerner2016raiders).
@@ 3.1. Data and Sampling
We produced a stratified frame seeded by Tranco, balanced by rank and vertical (news, retail, social, video) and bucketing regions as EU, US-CA, and Other for jurisdictional comparisons (pochat2019tranco). A headless browser, with new profile for each site, loads the homepage and privacy-policy page with geo/language hints to surface banners. For each snapshot, we retain: raw HTML, serialized consent-DOM subtree, full page screenshots of banners/modals, steps if they occur, and compact network log capturing request/cookie events with first/third-party flags. We clear storage between runs; the artefacts are content-hashed and time-stamped for de-duplication and reproducibility.
3.1.1. Banner Elicitation Settings.
@@ 3.2. Policy and UI Analysis
3.2.1. Policy clauses.
Each policy is parsed into clause candidates utilizing headings and sentence boundaries. Each clause, is sidled with similar embedded clauses across subsequent snapshots with a cost that incorporates semantic and edit similarity (reimers2019sbert; levenshtein1966). The clauses are labeled with a compact taxonomy (Purpose, Retention, Sharing, LegalBasis, Rights, Contact), and we consider the amount of clause churn as the share of clauses that substantially differ between months.
3.2.2. Consent patterns.
Consent UIs are classified by merging documented features (e.g., scrollable containers, toggle default selections, step indicators, calling text primary/secondary buttons/options, flags as visible or not) with screenshot markers. Weak rules confer training labels while a lighter classifier (a pooling image representation of documented features concatenated) predicts one of Scroll-Wall, Accordion, Multi-Step, Pre-ticked, or Reject-Hidden (ratner2017snorkel). A small, hand-labeled set aids in calibration and reporting precision/recall. We then calculate longitudinal prevalence and the row-normalized transition matrix (Table 3).
@@ 3.3. Metrics and Evaluation
3.3.1. Claim–UI alignment.
We pair policy claims (e.g., minimization, opt-in, easy reject) with necessary UI predicates (e.g., default-off, visible “Reject all”, steps-to-reject $\leq 2$) to obtain an alignment score $A\in[0,1]$ per snapshot. We summarize $A$ by region and vertical (Figure 1) and relate it to observed pattern shares.
3.3.2. Longitudinal analysis.
In the monthly series, we track the churn as well as the prevalence of patterns; we flag structural shifts with simple change point checks (killick2012changepoint). We summarize event responses as difference-in-differences with site and month fixed effects and use clustered inference (angrist2009mhe; wooldridge2010panel). Quality controls involve short-interval re-crawls for stability, and we report classifier metrics on the gold set.
3.4. Validation.
@@
3.4.1. Classifier generalization:
[…]
3.4.2. Sampling:
Tranco-stratified estimates remain stable under inverse-probability weights for banner surfacing by region and under rank weighting, with signs and relative magnitudes intact.
3.4.3. Banner-missing sensitivity.
Banner surfacing is asymmetric by region. We report IPW estimates using inverse probabilities of surfacing and a worst-case MNAR bound: assuming all missing EU (US–CA) banners take the 10th (90th) percentile of observed $A$ in that region, EU–US gaps shrink by at most 0.02; all signs and orderings remain.
3.4.4. Error propagation.
Propagating the observed confusion matrix via 1,000 bootstrap draws perturbs median $A$ by $\leq 0.02$ and DiD coefficients by $\leq 0.01$, leaving signs/orderings unchanged.
@@ 4.2. Effect Sizes and Robustness
4.2.1. Group gaps.
Median $A$ is higher in EU than US–CA by 0.09 (95% CI [0.07, 0.11]; Cliff’s $\delta=0.34$); News/Social exceed Retail. Visible “Reject all” associates with +0.12 $A$ [0.11, 0.13]; steps-to-reject $\leq 2$ adds +0.07 [0.05, 0.08]; default-off +0.05 [0.04, 0.06].
4.2.2. Event responses.
Post-enforcement we observe: “Reject all” +9.3 pp [6.8, 11.7], Pre-ticked -4.8 pp [-6.2, -3.4], and $A$ +0.04 [0.03, 0.05]; pre-trend tests are null (joint $p=0.41$).
4.2.3. Robustness.
Findings persist under rank weighting, excluding top–50 sites, alternate windows (±2/±3 months), EEA+UK bucketing, and vendor fixed effects (magnitudes within 0.01–0.02 of main).
@@ 4.3. Trends, Alignment, and Event Effects
We track monthly clause churn by taxonomy alongside consent-UI pattern prevalence; Figure 1 shows cross-sectional $A$ by region/vertical. We observe (hyp.): sustained churn in Purpose/Sharing in Q3–Q4, a gradual shift from Scroll-Wall toward Accordion, and a decline in Pre-ticked in EU buckets—consistent with template iteration and incremental CMP-layer adjustments.
@@
Table 1. Per-class performance on the 240-sample test set (bootstrap 95% CIs in parentheses).
@@ 4.4. Pattern Dynamics and Alignment
treating missing $u$ as unsatisfied. We use $w$ = {Reject-all visibility: 0.4; Default-off: 0.3; Steps-to-reject $\leq 2$: 0.2; Reopen/withdrawal affordance: 0.1}. Sensitivity to $w$ is reported below. An ablation indicates which predicates drive $A$ most: removing “Reject all” visibility causes the largest median drop, followed by default-off and steps-to-reject.
@@ 6. Conclusion
ConsentDiff presents an efficient, robust way to understand what’s in privacy policies and consent interfaces at the same time and on a web-scale. By combining clause-level policy diffs with a minimal DOM + vision classifier of consent interface patterns and claim-UI alignment score, we quantify temporal trends, transition dynamics, and jurisdictional differences without a user study. We find evidence for stable template implementations in policies; systematic differences in banners; and a higher degree of alignment where visible ”Reject all” and low steps-to-reject interfaces exist. We release code, schema, and a labeled subset of the data for replication and regulatory benchmarking. Future work will continue to vertically expand, build on multimodal classifiers with larger gold sets, and test the causal link between enforcement events and observed interface changes.

### v3

@@ Abstract.
Web privacy is experienced via two public artifacts: site utterances in policy texts, and the actions users are required to take during consent interfaces. In the extensive cross-section audits we’ve studied, there is a lack of longitudinal data detailing how these artifacts are changing together, and if interfaces are actually doing what they promise in policy. ConsentDiff provides that longitudinal view. We build a reproducible pipeline that snapshots sites every month, semantically aligns policy clauses to track clause-level churn, and classifies consent-UI patterns by pulling together DOM signals with cues provided by screenshots. We operationalize claim–UI alignment by mapping policy claims to observable UI predicates, connecting common policy claims to observable predicates, and enabling comparisons over time, regions, and verticals. Our measurements suggest continued policy churn, systematic changes to eliminate a higher-friction banner design, and significantly higher alignment where rejecting is visible and lower friction.
@@
journalyear: 2026copyright: ccconference: Extended Abstracts of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Extended Abstracts of the 2026 CHI Conference on Human Factors in Computing Systems (CHI EA ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772363.3798570isbn: 979-8-4007-2281-3/2026/04ccs: Security and privacy Human and societal aspects of security and privacyccs: Security and privacy Usability in security and privacyccs: Information systems Web miningccs: Information systems World Wide Web
@@ 1. Introduction
Online privacy is enacted through two public-facing artifacts: (i) privacy policies, which disclose purposes, retention, and sharing, and (ii) consent interfaces (CMP banners/modals), which implement user choice. Prior research has documented pervasive violations and dark-pattern designs in cookie banners and consent flows (Nouwens et al., 2020; Matte et al., 2020; Mathur et al., 2019). However, we still lack a longitudinal, web-scale view that jointly tracks what sites say in policy text and what users must do in consent interfaces, along with a concrete metric for whether these two artifacts align.
This is a problem for both end users and regulators. Policies are often updated without any notice, whether due to template refreshes, vendor switches, or jurisdiction-specific edits. Consent banners also change over time through A/B testing, CMP defaults, and implementation updates. Point-in-time audits can identify obvious violations, but they often miss how deployments evolve over time, which patterns persist, and whether policy claims keep pace with UI changes. At the same time, regulatory and industry guidance (e.g., recommendations for valid consent flows (European Data Protection Board, 2020; IAB Europe, 2020)) describe what compliant consent should look like, but they do not provide a reproducible way to measure whether common policy claims (e.g., “opt-in,” “easy reject”) correspond to the actual actions users must take.
We present ConsentDiff, a measurement framework that jointly tracks policy text and consent interfaces over time. Each month, we snapshot a stratified set of sites, semantically align policy clauses across versions, classify consent-UI patterns from DOM and screenshots, and compute a claim–UI alignment score that pairs policy statements with observable defaults and actions. Our work complements research on temporal drift in privacy recall which models how people misremember original sharing audience (Guo and Wei, 2026c); our focus is on the site-side artefacts, as well as noting their evolution.
We study how policy text and consent interfaces co-evolve at scale, guided by three research questions:
  1. RQ1:
    Claim–action consistency. When privacy policies make user-choice claims (e.g., easy rejection, opt-in, withdrawal), how often do the observed consent interfaces satisfy the corresponding UI predicates?
  2. RQ2:
    Longitudinal drift. How do policy clauses, consent-UI patterns, and claim–UI alignment change over time across regions and verticals?
  3. RQ3:
    Robust interpretation. How sensitive are longitudinal comparisons to imperfect banner surfacing and classifier error, and what diagnostics bound these risks?
@@ 2.1. Consent Interfaces and Dark Patterns
Large-scale audits have documented pervasive lack of compliance with and the use of dark-pattern designs in consent banners and CMP deployments, which illustrate how choices in interface can nudge users towards acceptance and obscure paths to rejection (Nouwens et al., 2020; Matte et al., 2020; Mathur et al., 2019; Degeling et al., 2019). Practice and policy guidance add further specificity to valid consent elements (e.g., unambiguous opt-in, symmetry of choices, easy withdrawal) and have codified vendor-facing frameworks that many sites adopt (European Data Protection Board, 2020; IAB Europe, 2020). Outside of consent dialogs, dark-pattern taxonomies and crawls have captured patterns of manipulative designs across e-commerce and platforms, including intrusion, default bias, and interference with interfaces (Mathur et al., 2019; Gray et al., 2021). Our focus on claim–UI alignment also relates to emerging HCI work on user-facing transparency cues in algorithmic systems, which similarly treats visible interface elements as measurable signals of what platforms reveal (or obscure) about system behavior and user control (Guo and Wei, 2026a). In our setting, we apply a similar interface-first lens to consent and privacy interfaces, but center whether user-choice claims in policy text are matched by observable, low-friction actions in the UI.
@@ 2.2. Automated Policy Analysis and Web-Scale Measurement
Automated tools segment and label privacy policies to enable designed query and compliance studies on a massive scale (Harkous et al., 2018; Lippi et al., 2019). Extending this area of work, we align policy clauses over time to measure clause-level churn, and to relate this churn to consent UI changes. There are research-oriented rankings that utilize methodology to reduce ”bias” in the top lists for our sampling and instrumentation (Pochat et al., 2019), which is supported with classic web-tracking studies on the importance of maintaining long term request/cookie logs and distinguishing between first/third parties, capturing screenshots indicating the UI states during data collection methods (Englehardt and Narayanan, 2016; Acar et al., 2014; Lerner et al., 2016). More broadly, our work sits in a line of web-facing measurement that treats public web artifacts as inputs to automated analysis pipelines and stresses the importance of robustness to real-world web variability and failure modes (Guo and Wei, 2026b).
@@ 3.1. Data and Sampling
We produced a stratified frame seeded by Tranco, balanced by rank and vertical (news, retail, social, video) and bucketing regions as EU, US-CA, and Other for jurisdictional comparisons (Pochat et al., 2019). A headless browser, with new profile for each site, loads the homepage and privacy-policy page with geo/language hints to surface banners. For each snapshot, we retain: raw HTML, serialized consent-DOM subtree, full page screenshots of banners/modals, steps if they occur, and compact network log capturing request/cookie events with first/third-party flags. We clear storage between runs; the artefacts are content-hashed and time-stamped for de-duplication and reproducibility.
@@ 3.2. Policy and UI Analysis
Policy clauses. Each policy is parsed into clause candidates utilizing headings and sentence boundaries. Each clause, is sidled with similar embedded clauses across subsequent snapshots with a cost that incorporates semantic and edit similarity (Reimers and Gurevych, 2019; Levenshtein, 1966). The clauses are labeled with a compact taxonomy (Purpose, Retention, Sharing, LegalBasis, Rights, Contact), and we consider the amount of clause churn as the share of clauses that substantially differ between months.
Policy clause extraction/alignment evaluation. In addition to the UI-pattern gold set, we evaluated policy processing on a hand-labeled subset of policy snapshots. We assessed (i) clause boundary quality against manual segmentation and (ii) month-to-month clause alignment accuracy on manually matched clause pairs spanning minor rewrites and template edits. Clause extraction was most reliable for heading-structured policies, while alignment errors concentrated in large template refreshes and pages with repeated boilerplate. We use these diagnostics to interpret clause-churn estimates conservatively and emphasize taxonomy-level trends over individual clause trajectories.
Consent patterns. Consent UIs are classified by merging documented features (e.g., scrollable containers, toggle default selections, step indicators, calling text primary/secondary buttons/options, flags as visible or not) with screenshot markers. Weak rules confer training labels while a lighter classifier, a pooling image representation of documented features concatenated, predicts one of Scroll-Wall, Accordion, Multi-Step, Pre-ticked, or Reject-Hidden (Ratner et al., 2017). A small, hand-labeled set aids in calibration and reporting precision/recall. We then calculate longitudinal prevalence and the row-normalized transition matrix (Table 3).
Gold set and agreement. We evaluate the UI-pattern classifier on a gold set of 240 snapshots, stratified by region and vertical (news, retail, social, video) with 20 samples per cell (12 cells total). A random subset of 96 items was double-coded using a shared codebook; inter-annotator agreement was Cohen’s $\kappa=\textbf{0.78}$, with disagreements resolved by adjudication.
@@ 3.3. Metrics and Evaluation
Claim–UI alignment. We pair policy claims with necessary UI predicates (e.g., default-off, visible “Reject all”, steps-to-reject $\leq 2$) to obtain an alignment score $A\in[0,1]$ per snapshot. We summarize $A$ by region and vertical (Figure 1) and relate it to observed pattern shares. Longitudinal analysis. In the monthly series, we track the churn as well as the prevalence of patterns; we flag structural shifts with simple change point checks (Killick et al., 2012). We summarize event responses as difference-in-differences with site and month fixed effects and use clustered inference (Angrist and Pischke, 2009; Wooldridge, 2010). Quality controls involve short-interval re-crawls for stability, and we report classifier metrics on the gold set.
3.4. Robustness Checks
We evaluate robustness along four axes: policy-processing reliability, UI-pattern classification error, banner non-surfacing, and sampling sensitivity. Manual checks on clause segmentation and temporal alignment indicate that policy-processing errors are concentrated in major template refreshes and repeated boilerplate; accordingly, we interpret clause churn conservatively and emphasize taxonomy-level trends.
@@ 3.4. Robustness Checks
Banner surfacing is asymmetric by region, so we report inverse-probability-weighted (IPW) estimates and a worst-case sensitivity bound. Banner non-surfacing is associated with script failures and late-loading UI; 27% of non-surfacing snapshots show JS console errors during elicitation, and IPW achieves covariate balance with max standardized mean difference 0.05. Under a worst-case MNAR bound, the EU–US gap shrinks by at most 0.02, with signs and relative orderings preserved.
Finally, Tranco-stratified estimates remain stable under rank weighting and inverse-probability weighting for banner surfacing by region, with effect directions and relative magnitudes unchanged.
@@ 4.2. Effect Sizes and Robustness
Group gaps. Median $A$ is higher in EU than US–CA by 0.09 (95% CI [0.07, 0.11]; Cliff’s $\delta=0.34$); News/Social exceed Retail. Visible “Reject all” associates with +0.12 $A$ [0.11, 0.13]; steps-to-reject $\leq 2$ adds +0.07 [0.05, 0.08]; default-off +0.05 [0.04, 0.06]. Event responses. Post-enforcement we observe: “Reject all” +9.3 pp [6.8, 11.7], Pre-ticked -4.8 pp [-6.2, -3.4], and $A$ +0.04 [0.03, 0.05]; pre-trend tests are null (joint $p=0.41$). Robustness. Findings persist under rank weighting, excluding top–50 sites, alternate windows (±2/±3 months), EEA+UK bucketing, and vendor fixed effects (magnitudes within 0.01–0.02 of main).
@@ 4.3. Trends, Alignment, and Event Effects
We track monthly clause churn by taxonomy alongside consent-UI pattern prevalence; Figure 1 shows cross-sectional $A$ by region/vertical. We observe: sustained churn in Purpose/Sharing in Q3–Q4, a gradual shift from Scroll-Wall toward Accordion, and a decline in Pre-ticked in EU buckets—consistent with template iteration and incremental CMP-layer adjustments.
@@
Table 1. Per-class UI-pattern classifier performance on the 240-sample test set.
@@ 4.4. Pattern Dynamics and Alignment
We treat undetected predicates as unsatisfied, yielding a conservative lower-bound estimate of alignment under imperfect observability. We use $w$ = {Reject-all visibility: 0.4; Default-off: 0.3; Steps-to-reject $\leq 2$: 0.2; Reopen/withdrawal affordance: 0.1}.
@@ 6. Conclusion
ConsentDiff presents an efficient, robust way to understand what’s in privacy policies and consent interfaces at the same time and on a web-scale. It enables reproducible, longitudinal audits and pre/post comparisons across regions and verticals. It can also evaluate CMP/UI interventions by tracking shifts in predicted patterns and alignment. By combining clause-level policy diffs with a minimal DOM + vision classifier of consent interface patterns and claim-UI alignment score, we quantify temporal trends, transition dynamics, and jurisdictional differences without a user study. We find evidence for stable template implementations in policies; systematic differences in banners; and a higher degree of alignment where visible ”Reject all” and low steps-to-reject interfaces exist. Future work will expand to other privacy-control surfaces beyond CMP banners, scale multimodal classifiers with larger stratified gold sets, and strengthen causal identification of event-aligned changes via richer event timing, alternative controls, and triangulation with independent evidence.